OpenStudy (anonymous):

One day a friend visited my house and asked me the ages of my three sons. I replied that the product of my son's ages is 36 and the sum of their ages is equal to my house number. My friend who has a sharp and logical mind, requested me for another hint. I then told him that my eldest son has blue eyes! Following this he quickly told me their ages

OpenStudy (anonymous):

Are we supposed the find their ages now? It's not possible from the data given right?

OpenStudy (bahrom7893):

x*y*z=36 x+y+z = # hmmm

OpenStudy (anonymous):

I request another hint too.

OpenStudy (bahrom7893):

no this is logic.. dang it my logic is not working..

OpenStudy (bahrom7893):

There's an eldest son..

OpenStudy (bahrom7893):

So 1*1*36.. that's far too fetched..

OpenStudy (bahrom7893):

36 = 9*4 So the ages could be 9 and 2 and 2

OpenStudy (bahrom7893):

* too far fetched

OpenStudy (bahrom7893):

36=6*3*2

OpenStudy (anonymous):

Aren't we supposed to know the house number?

OpenStudy (bahrom7893):

we are supposed to but i think the answer is 6, 3, 2

OpenStudy (anonymous):

i dnt knw i got only this info

OpenStudy (anonymous):

With this info there are multiple solutions.

OpenStudy (anonymous):

A. 1 year, 1 year and 36 years B. 1 year, 4 years and 9 years C. 2 years, 2 years and 9 years D. 3 years, 3 years and 4 years

OpenStudy (bahrom7893):

wait let's see..

OpenStudy (bahrom7893):

36 = 4*9 = 1*1*2*2*3*3

OpenStudy (anonymous):

anyone knw y did he give the hint blue eyes?

OpenStudy (anonymous):

So there is an eldest son.

OpenStudy (bahrom7893):

sophiya the hint is not blue eyes, the hint is ONE SON IS ELDEST..

OpenStudy (bahrom7893):

yea thomas was right.. hold on i think i got it..

OpenStudy (bahrom7893):

36 = 1*1*2*2*3*3, we need combinations of these: 1+1+(2*2*3*3) = 1+1+36=38 <<impossible

OpenStudy (bahrom7893):

1*1+2+(2*3*3) = 1 + 2 + (2*3*3) = 1 + 2 + 18 = 21 <<possible

OpenStudy (anonymous):

Wait, house number have an upper bound?

OpenStudy (anonymous):

hmm with this info hw can we get the answer?

OpenStudy (bahrom7893):

Keep doing this: 1*1+3+(2*2*3) = 1+3+12 = 16 1*1+(2*2)+(3*3)= 1+4+9 = 14 1*1+(2*3)+(2*3)= 1+6+6 = 13 2*1+2*1+(3*3) = 2+2+9 = 13 2*1+3*1+(2*3) = 2+3+6 = 11 3*1+3*1+(2*2) = 3+3+4 = 10

OpenStudy (bahrom7893):

Okay there we go those are the possible combinations of the sums.. See we have two of the same sum = 13 so ages are either: 1, 6, 6 or 2,2,9

OpenStudy (bahrom7893):

Now we use the fact that the OLDEST SON has blue eyes. Therefore there's one oldest son, not two.

OpenStudy (bahrom7893):

So we end up with: 2,2,9

OpenStudy (bahrom7893):

2,2,9 are the ages and 13 is the house number

OpenStudy (anonymous):

I don't see how it matters that two combinations have the same sum.

OpenStudy (anonymous):

atlast shall v end up with 2 2 9?

OpenStudy (anonymous):

hey bahroom thank u...

OpenStudy (bahrom7893):

Remember when the friend says that just the product is not enough: 1+1+(2*2*3*3) = 1+1+36=38 1*1+2+(2*3*3) = 1+2+18 = 21 1*1+3+(2*2*3) = 1+3+12 = 16 1*1+(2*2)+(3*3)= 1+4+9 = 14 1*1+(2*3)+(2*3)= 1+6+6 = 13 2*1+2*1+(3*3) = 2+2+9 = 13 2*1+3*1+(2*3) = 2+3+6 = 11 3*1+3*1+(2*2) = 3+3+4 = 10 See the friend is stuck with all these choices and i think the fact that there are two 13s is not a random coincidence..

OpenStudy (bahrom7893):

I was writing this out on paper when i noticed it..haha

OpenStudy (anonymous):

I get it. It's not yet possible to get the answer with just these combination. And that only happens when the house number is 13.

OpenStudy (anonymous):

i think v cant get answer with this info

OpenStudy (anonymous):

http://www.linuxquestions.org/questions/blog/dendron-600233/quick-puzzle-3981/ i got this que frm this link...

OpenStudy (anonymous):

hey guys at last anyone got d ans?

OpenStudy (anonymous):

hey how can it be 13 ok if there r two combinatios suming up to 13 but then how can we conclude tat it has to be tat

OpenStudy (anonymous):

i got it the friend knew the house number but was confused tat dere were two combinations resulting in 13 so he asked for the final clue i.e. the eldest son. yup tats true fr sure !!!!!

OpenStudy (anonymous):

The product of the ages of the three children is 36. Thus, their possible ages and the sum of their ages are: 1 year, 1 year, 36 years; Sum = 38 1 year, 3 years, 12 years; Sum = 16 1 year, 4 years, 9 years; Sum = 14 1 year, 6 years, 6 years; Sum = 13 2 years, 2 years, 9 years; Sum = 13 2 years, 3 years, 6 years; Sum = 11 3 years, 3 years, 4 years; Sum = 10 In this case, every combination of possible ages has a unique sum of ages - except for two cases, both of which have 13 as the sum. Since the friend could not figure out the ages after looking at the house number, the house number must be 13, because then, the ages could be either 1, 6, 6 or 2, 2, 9. When he was told that the eldest son has blue eyes, he immediately ruled out the set of 1 year, 6 years and 6 years as in this case, the eldest children have the same age. Thus, only one possibility remained, i.e., the ages of the children are 2 years, 2 years and 9 years.